# SVGFusion: Scalable Text-to-SVG Generation via Vector Space Diffusion

### v1

@@ 4.4 Ablation Study
[…]
VP-DiT S 12 384 6 1.4
VP-DiT B 12 768 12 5.6
VP-DiT L 24 1024 16 19.9
Table 2: Details of VP-DiT Models. We follow model configurations for the Small (S), Base (B) and Large (L) variants.
@@
Model Size. We apply a sequence of $\mathrm{N}$ DiT blocks, each operating at the hidden dimension size $d$. Following DiT [31], we use standard Transformer configs that jointly scale $\mathrm{N}$, $d$ and attention heads. Specifically, we use three configs: DiT-S, DiT-B, DiT-L. They cover a wide range of model sizes and flop allocations, from 1.4 to 19.9 Gflops, allowing us to gauge scaling performance. Table 2 gives details of the configs.
Figure 9: Comparison of Feature Extractors. We utilize pretrained DINOv2 [30] as our feature extractor to capture precise geometric features of objects at the pixel level. In contrast, [58, 65] trained ConvNet for feature extraction, which was less effective at pixel-level characterization and introduced significant background noise.
Comparison of Feature Extractors. As shown in Fig. 9, we utilize pretrained DINOv2 [30] as our feature extractor to capture precise geometric features of objects at the pixel level. In contrast, [58, 65] trained ConvNet for feature extraction, which was less effective at pixel-level characterization and introduced significant background noise.
We also conducted experiments to demonstrate the effectiveness of VP-VAE and Rendering Sequence Modeling. The results and analysis are provided in Supplementary Sec. A.
5 Conclusion
In this work, we have presented SVGFusion, a novel generative model for scalable vector graphics (SVG) creation. By leveraging a continuous vector space and eliminating the need for text-based discrete language models or extensive optimization processes, SVGFusion significantly advances the state-of-the-art in SVG generation. Our model efficiently transforms textual SVG code into compact vector embeddings and captures the underlying distributional properties of encoded vectors through a diffusion process. Additionally, the integration of vector primitive rendering sequences allows for hierarchical interpretation, closely mimicking human-like design patterns and mitigating common issues such as occlusion in complex graphic compositions. Extensive experiments have demonstrated the superior performance and scalability of SVGFusion. We believe our SVGFusion offers a promising new framework for SVG content creation across a wide range of applications.
@@ 5 Conclusion
This supplementary material is organized into several sections, each providing additional details and analyses related to our work on SVGFusion. Specifically, it covers the following aspects:
  • In Section A, we present the results of ablation studies to demonstrate the effects of VP-VAE and Rendering Sequence Modeling on the performance of SVGFusion.
@@
  • In Section B, we discuss the differences between the language model-based methods and our diffusion-based method, emphasizing how SVGFusion adopts an image-based paradigm to enhance SVG generation.
@@
  • In Section C, we demonstrate that SVGFusion supports a broader range of SVG primitives than existing methods, which enables it to capture the structural and visual characteristics of real-world SVGs more effectively.
  • In Section D, we provide additional details about the new dataset and data cleaning process.
  • In Section E, we present additional qualitative results of SVGFusion, demonstrating its capability to generate SVGs with high editability and visual quality.
Appendix A Additional Ablation Studies
Ablation on VP-VAE. In SVGFusion, we introduce the Visual-Pixel Fusion Autoencoder (VP-VAE) to learn a continuous latent space for SVGs. This module processes both SVG primitive parameters and corresponding rasterized images as inputs, aligning the SVG codes with visual features during the encoding phase. In Fig. S1, we compare results with and without the integration of visual features. We observe that incorporating visual features leads to higher generation quality.
Ablation on Rendering Sequence Modeling. The Rendering Sequence Modeling strategy in SVGFusion is designed to simulate the order in which humans create vector graphic. By processing batches of SVGs with incrementally added primitives—alongside both SVG codes and corresponding rasterizations—our model can learn the logic behind SVG creation. In Fig. S1, we compare the effects of employing this strategy versus omitting it. The results clearly show more logical outcomes, e.g., the relationships between the scoops of ice cream, when Rendering Sequence Modeling is applied. This improvement is due to the order of primitives in our generated SVGs being more aligned with human creation logic.
Figure S1: Effects of VP-VAE and Rendering Sequence Modeling. (a) vs. (c) demonstrates that without the Vector-Pixel Fusion (incorporating DINOv2 visual prior [30]), VP-VAE cannot accurately reconstruct or generate shapes. (a) vs. (c) indicates that employing Rendering Sequence Modeling results in more reasonable SVG outcomes, due to the order of primitives being better aligned with human creation logic.
Figure S2: Comparison of autoregressive language model prediction and diffusion-based generation of SVGFusion.
Appendix B Comparison with Language Model-based Methods
Instead, our SVGFusion utilizes a diffusion-based approach to synthesize shapes in parallel rather than sequentially. As demonstrated in the first row of Fig. S2, this method circumvents the inherent limitations of autoregressive methods, effectively reducing the accumulation of errors.
Appendix C More Primitive Types
SVG code comprises a suite of primitives and syntax rules. For example, the element-level primitive <rect> defines a rectangle shape with specified position, width, and height, and can be represented as <rect x="10" y="20" width="50" height="80"/>. However, given the multitude of SVG primitive types, it is challenging for Language Model-based methods to model these effectively due to the need for a complex data structure and extensive data. Consequently, some studies [3, 58, 59, 61, 50] simplify this task by focusing solely on one element-level primitive, <path>, and three command-level primitives: ”Move To” (M), ”Line To” (L), and ”Cubic Bézier” (C). Such simplification is lossy and thereby those methods fail to accurately represent real SVGs as designed by humans.
@@ Appendix C More Primitive Types
Primitive(element/command) Argument Explanation Example
@@
<path> Quadratic Bézier (Q) $\mu_{1},\nu_{1},\mu_{2},\nu_{2},\mu_{3},\nu_{3}$ Draws a quadratic Bézier curve with control points $(\mu_{1},\nu_{1})$ and endpoint $(\mu_{3},\nu_{3})$. [Uncaptioned image]
@@
Figure S4: SVG Icons Generated by SVGFusion.
@@
Figure S6: Description of the SVG cleaning process.
Appendix D SVG Dataset and Data Cleaning Process
SVG Dataset. The dataset consists of emoji and icon SVGs sourced from various sources, including Twemoji-Color-Font [54], Noto-Emoji [9], FluentUI-Emoji [25], SVG-Repo [49] and Reshot [37]. Twemoji-Color-Font, Noto-Emoji, and FluentUI-Emoji each contributes approximately 4,000 SVGs. An additional 30,000 SVGs are sourced from Reshot, while the majority—200,000 SVGs—are contributed by SVGRepo. These SVG resources are characterized by their high quality, rich color palettes, diverse designs, and human-designed nature, making them highly suitable for training and evaluation purposes.
Such a dataset enables us to design a model capable of leveraging prior knowledge of human-created SVGs to synthesize high-quality SVG outputs. However, SVG files collected from the web often contain noise, such as redundant data stemming from different designers’ workflows. To address this, we apply a preprocessing step to remove redundancy, enhance training efficiency, and ensure data consistency.
SVG Data Cleaning Process. The primary objective of the cleaning process is to perform lossless trimming of SVG files sourced from the web. Typically, more than half of the data in an SVG file is redundant and does not contribute to the rendering process. As illustrated in Figure S6, examples of rendering-irrelevant data include: 1) Temporary data generated by vector editing applications. 2) Suboptimal structural representations of SVG. 3) Unused or invisible graphical elements. These components are superfluous in finalized, published SVG files and can be safely removed to improve file efficiency.
Appendix E Additional Qualitative Results
Icon. As shwon in Fig. S4, the proposed method is also adept at synthesizing simple icon-style SVGs, which are typically composed of a limited number of primitives and are defined by their minimalistic yet expressive design characteristics.
Editability. Fig. S5 demonstrates the high editability of SVGs produced by SVGFusion. The clean and systematic structure of these SVGs facilitates efficient modification of primitive properties, including color attributes, enabling seamless customization.

### v2

@@ 4.4 Ablation Study
Model Size. We apply a sequence of $N_{b}$ DiT blocks, each operating at the hidden dimension size $d$. Following DiT [32], we use standard Transformer configs that jointly scale $N_{b}$, $d$ and attention heads. Specifically, we use three configs: VS-DiT S, VS-DiT B, VS-DiT L. They cover a wide range of model sizes and flop allocations, from 1.4 to 19.9 Gflops, allowing us to gauge scaling performance. The details of the configs are proviced in Table 2.
Model Layer $N_{b}$ Hidden size $d$ Heads Gflops
VS-DiT S 12 384 6 1.4
VS-DiT B 12 768 12 5.6
VS-DiT L 24 1024 16 19.9
Table 2: Details of VS-DiT Models. We follow model configurations for the Small (S), Base (B), and Large (L) variants.
Ablation on Vector-Pixel Fusion Encoding. In SVGFusion, we introduce the Visual-Pixel Fusion Autoencoder (VP-VAE) to learn a continuous latent space for SVGs. This module processes both SVG codes and their corresponding rendered images as inputs, aligning the SVG codes with visual features during the encoding phase. We evaluate the impact of integrating visual features by comparing results from model variants with and without this feature. As illustrated in Fig. 8, the variant incorporating Vector-Pixel Fusion Encoding demonstrates superior visual quality in the generated SVGs.
Furthermore, we conducted experiments to validate the effectiveness of the feature extractor, DINOv2, employed in our approach. Detailed results and analyses of these experiments are available in the Supplementary Sec. D.
Figure 8: Effects of VP-VAE and Rendering Sequence Modeling. (a) vs. (c) demonstrates that VP-VAE cannot accurately reconstruct or generate shapes without the Vector-Pixel Fusion (incorporating DINOv2 visual prior [31]). (b) vs. (c) indicates that employing Rendering Sequence Modeling results in more reasonable SVG outcomes, due to the order of primitives being better aligned with human creation logic.
Ablation on Rendering Sequence Modeling. The Rendering Sequence Modeling strategy in SVGFusion is designed to enhance the construction of generated SVGs, making them more editable. In Fig. 8, we evaluate the impact of employing this strategy compared to scenarios where it is omitted. The results clearly show that SVGs generated with Rendering Sequence Modeling exhibit higher visual quality and improved structural integrity.
In the first example, issues arise when shapes with lighter colors are occluded by shapes with darker colors, leading to a poorly represented money icon. In the second example, the creation order of the scoops is critical as their relational positioning greatly influences the visual coherence. These examples underscore the effectiveness of our Rendering Sequence Modeling strategy.
5 Conclusion & Discussion
In this work, we have introduced SVGFusion, a novel Text-to-SVG model that effectively scales to real-world vector graphics without relying on text-based discrete language models or prolonged optimization. By learning a continuous latent space through VP-VAE and generating structured SVG representations with VS-DiT, our approach captures the underlying creation logic of vector graphics, enabling high-quality and semantically coherent generation. To support this, we have curated a dataset called SVGX-Dataset. Extensive experiments validate the superiority of our SVGFusion over existing methods, demonstrating its potential as a scalable and efficient framework for SVG content creation.
Limitations. Due to the limited availability of publicly accessible data, the dataset collected in this work covers a restricted range of vector graphic types. Incorporating a broader diversity of vector graphics would not only expand the dataset size but also further explore the model’s scalability potential.
@@ 5 Conclusion & Discussion
This supplementary material provides additional details and analyses related to SVGFusion, organized as follows:
  • Appendix A: Implementation Details. We describe the implementation details of SVGFusion, including hyperparameter settings and training configurations.
  • Appendix B: Dataset Details. We provide an in-depth overview of our newly introduced dataset, covering data representation, preprocessing, and cleaning procedures.
  • Appendix C: VS-DiT Architecture. We present details of the Vector Space Diffusion Transformer (VS-DiT) architecture, explaining its components and design choices.
  • Appendix D Ablation Studies. We conduct ablation experiments to analyze the impact of different feature extractors on the performance of SVGFusion.
@@
  • Appendix E: Comparison with Language Model-Based Methods. We discuss the key differences between diffusion-based and language model-based SVG generation approaches, highlighting how SVGFusion leverages an image-based paradigm for improved SVG synthesis.
@@
  • Appendix F: SVG Primitive Support. We demonstrate that SVGFusion supports a broader range of SVG primitives than existing methods, enabling it to better capture the structural and visual diversity of real-world vector graphics.
  • Fig. S9: Additional Results. We provide qualitative results, including visualizations of the diffusion process and an extended set of generated SVGs, showcasing SVGFusion’s ability to black-and-white icon style vector graphics.
Appendix A Implementation Details
To ensure consistency across all SVG data, we adopted relative positional coordinates. The model parameters are initialized randomly and optimized using the AdamW optimizer (with $\beta_{1}=0.9$, $\beta_{2}=0.999$) at an initial learning rate of $3\times 10^{-4}$. The learning rate is warmed up over the first 2000 steps and then decayed to $1.5\times 10^{-5}$ following a cosine schedule. Additionally, we applied weight decay of 0.1 for regularization and constrain gradients by clipping their norms to a maximum value of 2.0. Besides, we normalized the input SVG embeddings into the $[-1,1]$ range to stabilize the training process. To further enhance the model’s ability to capture sequential dependencies, we integrate rotary position embeddings (RoPE) [48], a technique widely used in advanced large language models [30, 53]. RoPE effectively encodes positional relationships within sequences, allowing the model to better understand the temporal and structural progression of SVG creation. We utilized Transformers [55] as the fundamental building block for VP-VAE. Both the encoders and decoders are based on the Transformer architecture, consisting of 4 layers, a feed-forward dimension of 512. To investigate scaling trends, we trained VS-DiT models at three different sizes: 0.16B, 0.37B, and 0.76B parameters. We trained VP-VAE for 1000k steps using a total batch size of 128 across 2 A800 GPUs, requiring approximately two day. Subsequently, leveraging the trained VP-VAE, we trained VS-DiT for 500k steps with a batch size of 256 on 8 A800 GPUs, taking approximately four days. In the sampling phase, we used DDIM [43] to de-noise 100 steps; dpm-solver [24] also supports 20-step de-noising for more efficient sampling.
Figure S1: Samples from our collected SVG dataset. The dataset contains a diverse set of SVG illustrations, including outlined and filled designs, covering themes such as nature, objects, symbols, animals, and food.
Figure S2: Overview of our collected SVG dataset analysis. (a) displays the most frequently appearing words in the SVG dataset’s names and descriptions. Common words such as “arrow”, “circle”, and “left” suggest a predominance of directional and geometric elements in the dataset. (b) This word cloud visualizes the most common words found in the dataset. Larger words indicate higher frequency, highlighting key themes such as geometric shapes (“circle”, “square”), directional terms (“left”, “right”), and common UI elements (“file”, “folder”). (c) compares the number of different SVG elements (e.g., <path>, <rect>, <circle>) before and after data cleaning. The reduction in element count suggests significant optimization, potentially removing redundant or unnecessary elements.
Figure S3: Overview of the SVG cleaning process. The pipeline illustrates the transformation of an original SVG file into a preprocessed version through an SVG cleaning pipeline. The pipeline involves several steps: removing XML declarations, comments, titles, and descriptions; processing and eliminating <defs> and <class> attributes; removing unused <id> tags and global <g> tags; converting absolute coordinates to relative; setting path coordinate precision to 2; and resizing the canvas to $128\times 128$ pixels. The cleaned SVG results in a more optimized and compact representation while preserving the visual integrity of the image.
Appendix B SVG Dataset Collection, Representation, Cleaning and Analysis
SVG Data Collection. We collected a large-scale SVG dataset for model training and evaluation, which includes 240k high-quality SVGs with emoji and icon style. The dataset consists of emoji and icon SVGs sourced from various sources, including Twemoji-Color-Font [54], Noto-Emoji [10], FluentUI-Emoji [27], SVG-Repo [62] and Reshot [61]. Twemoji-Color-Font, Noto-Emoji, and FluentUI-Emoji each contribute approximately 4,000 SVGs. An additional 30,000 SVGs are sourced from Reshot, while the majority—200,000 SVGs—are contributed by SVGRepo.
SVG Data Representation In terms of visual complexity, our collected dataset includes primitives that vary from simple to intricate. Regarding primitive representation, the dataset includes both Bézier curves expressed using <path> and basic shapes represented by <circle> and <rect>, etc. As illustrated in Fig. S1, in terms of color representation, the dataset contains both black-and-white graphics and vibrant, harmonious color images. Additionally, in terms of semantics, it covers a wide range of subjects, including people, animals, objects, and symbols. Such a dataset enables us to design a model capable of leveraging prior knowledge of human-created SVGs to synthesize high-quality SVG outputs.
SVG Data Cleaning Pipeline. However, SVG files collected from the web often contain noise, such as redundant data stemming from different designers’ workflows, and using it directly for learning can hinder the model’s representation capabilities. As illustrated in Fig. S3, about half of the data in an SVG file is redundant for visual rendering, including (1) temporary data from vector editing applications, (2) non-optimal structural representations, and (3) unused or invisible graphic elements. These components are superfluous in finalized, published SVG files and can be safely removed to improve file efficiency. We propose a preprocessing pipeline that losslessly reduces the size of SVG files by eliminating unnecessary XML declarations, comments, titles, and descriptions, as well as removing unused SVG tags. Additionally, we convert absolute coordinates to relative ones, set the precision of path coordinates to two decimal places, and resize the canvas to $128\times 128$.
@@ Appendix B SVG Dataset Collection, Representation, Cleaning and Analysis
Figure S5: The structure of the VS-DiT Block. The architecture of the VS-DiT Block, where a multi-head cross-attention module is incorporated into each block to inject textual conditions. The text prompt is first processed through a CLIP text encoder and a linear transformation before being integrated into the model. The block consists of alternating multi-head self-attention and feed-forward layers, with scale and shift operations applied at multiple stages. Additionally, time-dependent conditioning is introduced through an MLP module. The block is repeated $N_{b}$ times to refine the latent representation.
Figure S6: Comparison of Feature Extractors. We employ pretrained DINOv2 [31] as our feature extractor, as it provides a more precise spatial-geometric representation of objects without interference from background noise. In contrast, previous methods [58, 69] relied on ConvNet-based feature extraction, which introduced additional background noise, ultimately degrading the quality of geometric representations.
SVG Dataset Analysis. As illustrated in Fig. S2, our dataset analysis focuses on three key aspects: (a) the 40 most frequently occurring words in the names of the SVG dataset, (b) a word cloud representing the most common words found in SVG names and descriptions, and (c) a comparative analysis of the dataset before and after the cleaning process.
Appendix C VS-DiT Block Architecture
As illustrated in Fig. S5, it is positioned between the self-attention layer and the feed-forward layer so that the model can flexibly interact with the text embedding extracted from the language model. Our architecture supports flexible stacking of $N_{b}$ VS-DiT blocks, enabling models with varied parameter scales that allow SVGFusion to adaptively expand with increasing data volume, thus enhancing scalability with data growth.
Appendix D Additional Ablation Studies
Comparison of Feature Extractors. As shown in Fig. S6, we utilize pretrained DINOv2 [31] as our feature extractor to capture precise geometric features of objects at the pixel level. In contrast, [58, 69] trained ConvNet for feature extraction, which was less effective at pixel-level characterization and introduced significant background noise.
Appendix E Comparison with Language Model-based Methods
SVG Primitives (Element/Command) Argument Explanation Example
@@ Appendix E Comparison with Language Model-based Methods
<path> Quadratic Bézier (Q) $\mu_{1},\nu_{1},\mu_{3},\nu_{3}$ Draws a quadratic Bézier curve with control points $(\mu_{1},\nu_{1})$ and endpoint $(\mu_{3},\nu_{3})$. [Uncaptioned image]
@@
Instead, our SVGFusion employs a diffusion-based approach to synthesize shapes in parallel rather than sequentially. As shown in the second row of Fig. S7, the process begins with random, chaotic strokes, and the SVG elements, such as vector coordinates, are progressively denoised at each timestep. This method overcomes the inherent limitations of autoregressive techniques by mitigating error accumulation. Over successive iterations, the correct content gradually emerges while preserving the visual coherence of the generated SVG.
Figure S7: Comparison of autoregressive language model prediction and diffusion-based generation of SVGFusion. This comparison highlights the differences between two approaches for SVG generation: autoregressive language models and diffusion-based methods. Autoregressive models generate vector graphics sequentially, predicting one token at a time based on prior outputs, which can lead to error accumulation and limited global coherence. In contrast, SVGFusion leverages diffusion-based generation, which operates in a continuous vector space, enabling more holistic and globally consistent SVG synthesis while mitigating the limitations of discrete token-based generation.
Appendix F More Primitive Types
SVG code comprises a suite of primitives and syntax rules. For example, the element-level primitive <rect> defines a rectangle shape with specified position, width, and height, and can be represented as <rect x="10" y="20" width="50" height="80"/>. However, given the multitude of SVG primitive types, it is challenging for Language Model-based methods to model these effectively due to the need for a complex data structure and extensive data. Consequently, some studies [4, 58, 59, 63, 49] simplify this task by focusing solely on one element-level primitive, <path>, and three command-level primitives: “Move To” (M), “Line To” (L), and “Cubic Bézier” (C). Such simplification is lossy and thereby those methods fail to accurately represent real SVGs as designed by humans.
Appendix G Additional Qualitative Results
Figure S8: A Set of black-and-white SVG Icons Generated Using SVGFusion, illustrating various symbols representing weather, time, shapes, and miscellaneous objects.
@@ Appendix G Additional Qualitative Results
Icon. As shown in Fig. S8, the proposed method is also adept at synthesizing simple black-and-white icon-style SVGs, which are typically composed of a limited number of primitives and are defined by their minimalistic yet expressive design characteristics.
Editability. Figure S9 demonstrates the high editability of SVGs produced by SVGFusion. The clean and systematic structure of these SVGs facilitates efficient modification of primitive properties, including color attributes, enabling seamless customization.